OpenStudy (anonymous):

limx→0 3x+1−cos2(x)   sin(x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ 3x+(1-\cos^2(x) }{ \sin(x) }\]=?

OpenStudy (anonymous):

@Preetha

OpenStudy (cggurumanjunath):